One hundred of the voters in a town are willing to pay $100 each to support a public green space, which will cost $10,000 to build and maintain. One hundred and fifty voters in the same town do not value the public green space. What is the social marginal benefit of the public green space? $0. $100. $5,000. $10,000.
Answer #1

An individual is Willing to maximum amount of money for public green space is equal to marginal benefit from that public green space.

So each 100 individuals are willing to pay 100$ .so total benefit from that public green space=100*100=10,000

150$ doesn't value green space,means they get zero marginal benefit from green space.

So their total benefit=150*0=0

The society marginal benefit=100*100+150*0=10,000

So social marginal benefit is 10,000.

Option D is correct

Note: The net social marginal benefit will equal to =10,000- cost of green space =10,000-10,000=0